OpenStudy (anonymous):

A rectangular garden has a width of 5 feet and a length of 10 feet. If an equal amount is added to both the width and the length, the area is increased to 84 square feet. What is this amount?

OpenStudy (frostbite):

Let me see if i understand right? T = area, w = width, l = length, k = the constant amount. T = (w+k)*(l+k) 84 = (5 + k) * (10 + k) = k^2 + 15k + 50 0 = k^2+15k - 34 This is a quadratic equation so: D = 15^2-4*1*(-34)=361 k = (-15 (+-) sqrt(361))/(1*2) k = (-15 (+-) 19)/2 k = 2 V k = -17 then again it is hard to make a negative garden so 2 most be the answer. (if the premises i made in the start are correct)

OpenStudy (anonymous):

thanks that sounds right